# Tradeoffs of Diagonal Fisher Information Matrix Estimators

### v2

@@ Abstract
The Fisher information matrix characterizes the local geometry in the parameter space of neural networks. It elucidates insightful theories and useful tools to understand and optimize neural networks. Given its high computational cost, practitioners often use random estimators and evaluate only the diagonal entries. We examine two such estimators, whose accuracy and sample complexity depend on their associated variances. We derive bounds of the variances and instantiate them in regression and classification networks. We navigate trade-offs of both estimators based on analytical and numerical studies. We find that the variance quantities depend on the non-linearity w.r.t. different parameter groups and should not be neglected when estimating the Fisher information.
@@
This is the Supplementary Material to Paper "Tradeoffs of Diagonal Fisher Information Matrix Estimators". To differentiate with the numberings in the main file, the numbering of Theorems is letter-based (A, B, …).
@@ 1 Settings
In the parameter space of neural networks (NNs), i.e. the neuromanifold Amari (2016), the network weights and biases play the role of a coordinate system and the local metric tensor can be described by the Fisher Information Matrix (FIM). As a result, empirical estimation of the FIM helps reveal the geometry of the loss landscape during parameter learning and the intrinsic structure of the neuromanifold. Utilizing these insights has lead to efficient optimization algorithms, e.g., the natural gradient Amari (2016) and Adam Kingma & Ba (2015).
A NN with inputs $\bm{x}$ and stochastic outputs $\bm{y}$ can be specified by the conditional probability distribution $p(\bm{y}\,|\,\bm{x};\bm{\theta})$, where $\bm{\theta}$ is the NN weights and biases. This paper considers the general parametric form
@@
where $\bm{h}_{\bm{\theta}}:\Re^{I}\rightarrow\Re^{T}$ maps $I$-dimensional inputs $\bm{x}$ to $T$-dimensional exponential family parameters, $\bm{t}(\bm{y})$ is a vector of sufficient statistics, $\pi(\bm{y})$ is a base measure, and $F(\cdot)$ is the log-partition function (normalizing the exponential). For example, if $\bm{y}$ denotes class labels and $\bm{t}(\bm{y})$ maps to its corresponding one-hot vectors, then Eq. 1 is associated with a multi-class classification network.
@@
By assuming that the marginal data distribution $q(\bm{x})$ is parameter-free, we get parametric models $p(\bm{x},\bm{y};\bm{\theta})=q(\bm{x})p(\bm{y}\,|\,\bm{x};\bm{\theta})$ of the joint distribution. By definition, the (joint) FIM has the form $\mathcal{I}(\bm{\theta})\vcentcolon=\mathop{\mathbb{E}}_{q(\bm{x})}\left[\mathcal{I}(\bm{\theta}\,|\,\bm{x})\right]$, where
@@
$\mathcal{I}(\bm{\theta}\,|\,\bm{x})$ :$\vcentcolon=\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\frac{\partial\log p(\bm{y}\,|\,\bm{x};\bm{\theta})}{\partial\bm{\theta}}\frac{\partial\log p(\bm{y}\,|\,\bm{x};\bm{\theta})}{\partial\bm{\theta}^{\top}}\right]{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\overset{(*)}{=}}-\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\frac{\partial^{2}\log p(\bm{y}\,|\,\bm{x};\bm{\theta})}{\partial\bm{\theta}\partial\bm{\theta}^{\top}}\right]$ (2)
@@
is the ‘conditional FIM’ and $\mathop{\mathbb{E}}$ denotes the expectation. The second equality (*) holds if the activation functions are in $C^{2}(\Re)$ (i.e., $\bm{h}_{\bm{\theta}}$ is a sufficiently smooth NN). $\mathcal{I}(\bm{\theta}\,|\,\bm{x})$ does not have this equivalent expression (*) for NNs with ReLU activation functions Soen & Sun (2021). Both $\mathcal{I}(\bm{\theta})$ and $\mathcal{I}(\bm{\theta}\,|\,\bm{x})$ give a $\dim(\bm{\theta})\times\dim(\bm{\theta})$ positive semi-definite (PSD) matrix. The distinction in notation is for emphasizing that the joint FIM (depending only on $\bm{\theta}$) is simply the average over individual conditional FIMs (depending on both $\bm{\theta}$ and $\bm{x}$).
@@
In practice, the FIM is typically computationally expensive and needs to be estimated. Given $q(\bm{x})$ and a NN with weights and bias $\bm{\theta}$ parameterizing $p(\bm{y}\,|\,\bm{x};\bm{\theta})$, as per Eq. 1, we consider two commonly used estimators of the FIM Guo & Spall (2019); Soen & Sun (2021) given by
@@
^$\hat{\mathcal{I}}_{1}(\bm{\theta})$ :$\vcentcolon=\frac{1}{N}\sum_{k=1}^{N}\left[\frac{\partial\log p(\bm{y}_{k}\,|\,\bm{x}_{k})}{\partial\bm{\theta}}\frac{\partial\log p(\bm{y}_{k}\,|\,\bm{x}_{k})}{\partial\bm{\theta}^{\top}}\right],$ (3)
^$\hat{\mathcal{I}}_{2}(\bm{\theta})$ :$\vcentcolon=\frac{1}{N}\sum_{k=1}^{N}\left[-\frac{\partial^{2}\log p(\bm{y}_{k}\,|\,\bm{x}_{k})}{\partial\bm{\theta}\partial\bm{\theta}^{\top}}\right],$ (4)
@@
where we shorthand $p(\bm{y}_{k}\,|\,\bm{x}_{k})\vcentcolon=p(\bm{y}_{k}\,|\,\bm{x}_{k};\bm{\theta})$ and $(\bm{x}_{1},\bm{y}_{1}),\ldots,(\bm{x}_{N},\bm{y}_{N})$ are i.i.d. sampled from $p(\bm{x},\bm{y};\bm{\theta})$. A conditional variant of the above estimators, denoted as ^$\hat{\mathcal{I}}_{1}(\bm{\theta}\,|\,\bm{x})$ and ^$\hat{\mathcal{I}}_{2}(\bm{\theta}\,|\,\bm{x})$, can be defined by fixing $\bm{x}=\bm{x}_{1}=\cdots=\bm{x}_{N}$ and sampling $\bm{y}_{1},\ldots,\bm{y}_{N}$ from $p(\bm{y}\,|\,\bm{x};\bm{\theta})$ in Eqs. 3 and 4 — details omitted for brevity.
@@
Both estimators are random matrices with the same shape as $\mathcal{I}(\bm{\theta})$. By Eq. 2, their means are exactly the FIM and thus they are unbiased [^1] Only true for ^$\hat{\mathcal{I}}_{2}(\bm{\theta})$ if activation functions are in $C^{2}(\Re)$.. For the setting presented in Eq. 1, the variances of the estimators Eqs. 3 and 4 can be in closed form and upper-bounded Soen & Sun (2021). This provides an important tool in characterizing the accuracy of the estimators to gain insights on where/when different estimators of the FIM should be used. Despite this, neither the variance nor the bounds can be computed efficiently in practice due to the huge dimensionality of $\bm{\theta}$ in deep NNs.
This work focuses on estimating the diagonal entries of the FIM and their associated variance. Our main results — including the estimators of the FIM, their variances, and the bounds of the variances — can be computed efficiently through automatic differentiation and are therefore useful in practice. Our analytical results reveal how the moments of the output exponential family and how the gradients of the neural network in Eq. 1 affects the estimation of the FIM. We discover a general decomposition of the variance of the FIM corresponding respectively to the sampling of $\bm{x}$ and $\bm{y}$. We further give an estimation formula that can trade-off between the bias and variance in estimating the FIM and discuss its meaning in deep learning.
@@ 2 Related Work
Prior efforts aim to analyze the structure of the FIM of neural networks with randomized weights Amari et al. (2019). This body of work hinges on utilizing tools from random matrix theory and spectral analysis, characterizing the behavior and statistics of the FIM Pennington & Worah (2018). One insight is that randomly weighted neural networks have FIMs with a majority of eigenvalues close to zero; with the other eigenvalues taking large values Karakida et al. (2019; 2021). In our work, the randomness stems from sampling from data distributions $p(\bm{y}\,|\,\bm{x}_{k})$. Despite this difference, we note that our settings intersect at one point of the learning process of a neural network: the initial training iterations. Indeed, the random matrices examined in these works provide insights about the initialization of neural networks Karakida et al. (2019). Issues for utilizing the empirical FIM to approximate the FIM have been highlighted Pascanu & Bengio (2014); Martens (2020). For example, estimators of the FIM do not in general capture any second-order information about the log-likelihood Kunstner et al. (2020).
Various structural FIM approximations have been considered to examine trade-offs between second-order information and computation Karakida & Osawa (2020). One can consider the unit-wise FIM Ollivier (2015); Le Roux et al. (2007); Sun & Nielsen (2017); Amari et al. (2019), where a block-diagonal approximation of the FIM is taken to capture curvature information of intra-neuron parameter groups. Alternatively, one can study the block-diagonal layer-wise FIM which considers the parameter interaction within each layer Kurita (1994); Martens et al. (2010); Pascanu & Bengio (2014); Martens & Grosse (2015); Heskes (2000); Ren & Goldfarb (2021). Structurally, the diagonal FIM is a subset of the unit-wise FIM, which in turn is a subset of the layer-wise FIM. In the literature, a class of FIM estimators denoted as the quasi-diagonal FIM actually approximates the unit-wise FIM Ollivier (2015); Amari et al. (2019). Alternative to the FIM is the Generalized Gauss-Newton (GGN) matrix — a Hessian approximator — which was originally motivated through the squared loss for non-linear models Martens (2020). The GGN is shown to be equivalent to the FIM when a loss function is taken to be the empirical expectation of the negative log-likelihood of Eq. 1 Heskes (2000); Pascanu & Bengio (2014); Martens (2020).
The FIM estimators for NNs follow the same principle as Monte Carlo (MC) information geometry Nielsen & Hadjeres (2019) where information geometric quantities are approximated via MC estimation.
@@ 3 Variance of Diagonal FIM Estimators
In our notations, all vectors such as $\bm{x}$, $\bm{y}$, and $\bm{\theta}$ are column vectors. We use $k$ to index random samples $\bm{x}$ and $\bm{y}$ and use $i$ and $j$ to index the neural network weights and biases $\bm{\theta}$. We shorthand $\bm{h}:=\bm{h}_{\bm{\theta}}$, $p(\bm{y}\,|\,\bm{x})\vcentcolon=p(\bm{y}\,|\,\bm{x};\bm{\theta})$, and $p(\bm{y},\bm{x})\vcentcolon=p(\bm{y},\bm{x};\bm{\theta})$ whenever the parameters $\bm{\theta}$ is clear from context. To be consistent, we use ‘$\,|\,\bm{x}$ conditioning’ to distinguish between jointly calculated values versus conditioned values with fixed $\bm{x}$. By default, the derivatives are w.r.t. $\bm{\theta}$. For example, $\partial_{i}\bm{h}\vcentcolon=\partial\bm{h}/\partial\theta_{i}$ and $\partial_{i}^{2}\bm{h}\vcentcolon=\partial^{2}\bm{h}/\partial\theta_{i}^{2}$. We adopt Einstein notation to express tensor summations, so that an index appearing as both a subscript and a superscript in the same term indicates a summation. For example, $x^{a}y_{a}$ denotes $\sum_{a}x^{a}y_{a}$. Whenever convenient, we use a mix of standard $\Sigma$-sum and Einstein notation.
@@
Based on the parametric form of the model in Eq. 1, the diagonal entries of the FIM estimators in Eqs. 3 and 4 can be written as follows [^2] Derivations of our formal results are provided in the appendix.:
@@
^$\hat{\mathcal{I}}_{1}(\theta_{i})$ $=\frac{1}{N}\sum_{k=1}^{N}\left(\frac{\partial F(\bm{h}(\bm{x}_{k}))}{\partial\theta_{i}}-\frac{\partial\bm{h}^{a}(\bm{x}_{k})}{\partial\theta_{i}}\cdot\bm{t}_{a}(\bm{y}_{k})\right)^{2},$ (5)
Clearly ^$\hat{\mathcal{I}}_{1}(\theta_{i})\geq 0$, while there is no such guarantee for ^$\hat{\mathcal{I}}_{2}(\theta_{i})$ which can be negative.
Our results will be expressed in terms of the (central) moments of $\bm{t}(\bm{y})$:
$\bm{\eta}_{a}(\bm{x})$ :$\vcentcolon=\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x})}[\bm{t}_{a}(\bm{y})],$ (7)
$\mathcal{I}(\bm{h}\,|\,\bm{x})$ :$\vcentcolon=\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x})}[(\bm{t}(\bm{y})-\bm{\eta}(\bm{x}))(\bm{t}(\bm{y})-\bm{\eta}(\bm{x}))^{\top}],$ (8)
$\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})$ :$\vcentcolon=\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x})}\left[(\bm{t}(\bm{y})-\bm{\eta}(\bm{x}))\otimes(\bm{t}(\bm{y})-\bm{\eta}(\bm{x}))\otimes\;(\bm{t}(\bm{y})-\bm{\eta}(\bm{x}))\otimes(\bm{t}(\bm{y})-\bm{\eta}(\bm{x}))\right],$ (9)
where $\otimes$” denotes the tensor product. We denote the covariance of $\bm{t}$ w.r.t. to $p(\bm{y}\,|\,\bm{x})$ as $\mathrm{Cov}^{p}(\bm{t}\,|\,\bm{x})$ — noting that Eq. 8 is simply $\mathcal{I}(\bm{h}\,|\,\bm{x})=\mathrm{Cov}^{p}(\bm{t}\,|\,\bm{x})$. The 4D tensor $\mathcal{K}^{p}(\bm{t}\,|\,\bm{x})$ denotes the $4^{\mathrm{th}}$ central moment of $\bm{t}(\bm{y})$ w.r.t. $p(\bm{y}\,|\,\bm{x})$. These (central) moments correspond to the cummulants of $\bm{t}(\bm{y})$, i.e. the derivatives of $F$ w.r.t. the natural parameters $\bm{h}(\bm{x})$ of the exponential family. Therefore, the derivatives of $F$ in Eqs. 5 and 6 can further be written in terms of $\bm{\eta}(\bm{x})$ and $\mathcal{I}(\bm{h}\,|\,\bm{x})$ following the chain rule.
In practice, both estimators can be computed via automatic differentiation Paszke et al. (2019); Dangel et al. (2019). In terms of complexity, by restricting our estimates of the FIM to just its diagonal elements, we only need to calculate $\mathcal{O}(\dim(\bm{\theta}))$ elements (originally $\mathcal{O}(\dim(\bm{\theta})\times\dim(\bm{\theta}))$ for the full FIM). Although the log-partition function for general exponential family distributions can be complicated, for the ones used in neural networks (determined by the loss functions used in optimization) Soen & Sun (2021) the log-partition function $F$ can be in closed-form; and thus the cummulants in Eqs. 7 and 8 can be calculated efficiently.
Indeed, the primary cost of the estimators comes from evaluating the gradient information of the neural network, given by $\partial\bm{h}(\bm{x})/\partial\theta_{i}$ and $\partial^{2}\bm{h}(\bm{x})/\partial^{2}\theta_{i}$. The former can be calculated easily. The latter is costly even when restricted to the diagonal elements of the FIM. With the Hessian’s quadratic complexity, in practice approximations are used to reduce the computational overhead Becker et al. (1988); Yao et al. (2020; 2021); Elsayed & Mahmood (2022). In this case, additional error and (potentially) variance may be introduced as a result of the Hessian approximation. It should however be noted that if we only wish to calculate the diagonal entries of the FIM for layers close the output, the cost of computing the Hessian can still be manageable. Indeed, to calculate the Hessians required for the FIM, we require all Hessian matrices (w.r.t. layer inputs) for layers closer to the output by the chain rule. As such, part of the cost can be circumvented by only considering the outermost layers. Despite this, one will still pay a cost in memory which quadratically scales with the number of parameters for non-linear activation functions (Dangel et al., 2019).
The high cost associated with the Hessian computation does not justify that ^$\hat{\mathcal{I}}_{1}$ should always be used. Such a cost can be outweighed by the FIM estimator’s variance in particular settings (choices of loss functions) Soen & Sun (2021). Indeed, computation is not the only determining factor for the usefulness of the estimator. This is especially true when the FIM can be used in an offline setting — where the Hessian’s cost can be tolerated — to study the singular structure of the neuromanifold Amari et al. (2018); Sun & Nielsen (2019), the curvature of the loss Efron (2018), to quantify model sensitivity Nickl et al. (2023), and to evaluate the quality of the local optimum Karakida et al. (2019; 2021), etc.
To study the quality of ^$\hat{\mathcal{I}}_{1}(\bm{\theta})$ and ^$\hat{\mathcal{I}}_{2}(\bm{\theta})$, it is natural to examine the variance of the estimators Soen & Sun (2021):

### v3

# Trade-Offs of Diagonal Fisher Information Matrix Estimators

@@ Abstract
The Fisher information matrix can be used to characterize the local geometry of the parameter space of neural networks. It elucidates insightful theories and useful tools to understand and optimize neural networks. Given its high computational cost, practitioners often use random estimators and evaluate only the diagonal entries. We examine two popular estimators whose accuracy and sample complexity depend on their associated variances. We derive bounds of the variances and instantiate them in neural networks for regression and classification. We navigate trade-offs for both estimators based on analytical and numerical studies. We find that the variance quantities depend on the non-linearity w.r.t. different parameter groups and should not be neglected when estimating the Fisher information.
@@
This is the Supplementary Material to Paper "Trade-Offs of Diagonal Fisher Information Matrix Estimators
The current article is a digital reprint of Soen and Sun [2024].". To differentiate with the numberings in the main file, the numbering of Theorems is letter-based (A, B, …).
@@ 1 Settings
In the parameter space of neural networks (NNs), i.e. the neuromanifold Amari (2016), the network weights and biases play the role of a coordinate system and the local metric tensor can be described by the Fisher Information Matrix (FIM). As a result, empirical estimation of the FIM helps reveal the geometry of the loss landscape and the intrinsic structure of the neuromanifold. Utilizing these insights has lead to efficient optimization algorithms, e.g., the natural gradient Amari (2016) and Adam Kingma and Ba (2015).
A NN with inputs $\bm{x}$ and stochastic outputs $\bm{y}$ can be specified by a conditional p.d.f. $p(\bm{y}\,|\,\bm{x};\bm{\theta})$, where $\bm{\theta}$ is the NN’s weights and biases. This paper considers the general parametric form
@@
where $\bm{h}_{\bm{\theta}}\colon\Re^{I}\rightarrow\Re^{T}$ maps $I$-dimensional inputs $\bm{x}$ to $T$-dimensional exponential family parameters, $\bm{t}(\bm{y})$ is a vector of sufficient statistics, $\pi(\bm{y})$ is a base measure, and $F(\cdot)$ is the log-partition function (normalizing the exponential). For example, if $\bm{y}$ denotes class labels and $\bm{t}(\bm{y})$ maps to its corresponding one-hot vectors, then Eq. 1 is associated with a multi-class classification network.
@@
Assuming that the marginal distribution $q(\bm{x})$ is parameter-free, we define parametric joint distributions $p(\bm{x},\bm{y};\bm{\theta})=q(\bm{x})p(\bm{y}\,|\,\bm{x};\bm{\theta})$. The (joint) FIM is defined as $\mathcal{I}(\bm{\theta})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\mathop{\mathbb{E}}_{q(\bm{x})}\left[\mathcal{I}(\bm{\theta}\,|\,\bm{x})\right]$, where
@@
$\mathcal{I}(\bm{\theta}\,|\,\bm{x})$ =.$\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\frac{\partial\log p(\bm{y}\,|\,\bm{x};\bm{\theta})}{\partial\bm{\theta}}\frac{\partial\log p(\bm{y}\,|\,\bm{x};\bm{\theta})}{\partial\bm{\theta}^{\top}}\right]{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\overset{(*)}{=}}-\mathop{\mathbb{E}}_{p(\bm{y}\,|\,\bm{x};\bm{\theta})}\left[\frac{\partial^{2}\log p(\bm{y}\,|\,\bm{x};\bm{\theta})}{\partial\bm{\theta}\partial\bm{\theta}^{\top}}\right]$ (2)
@@
is the ‘conditional FIM’. The second equality (*) holds if $\bm{h}_{\bm{\theta}}$’s activation functions are in $C^{2}(\Re)$ (i.e., $\bm{h}_{\bm{\theta}}$ is a sufficiently smooth NN). $\mathcal{I}(\bm{\theta}\,|\,\bm{x})$ does not have this equivalent expression (*) for NNs with ReLU activation functions Soen and Sun (2021). Both $\mathcal{I}(\bm{\theta})$ and $\mathcal{I}(\bm{\theta}\,|\,\bm{x})$ define $\dim(\bm{\theta})\times\dim(\bm{\theta})$ positive semi-definite (PSD) matrices. The distinction in notation is to emphasize that the joint FIM $\mathcal{I}(\bm{\theta})$ (depending only on $\bm{\theta}$) is simply the average over individual conditional FIMs $\mathcal{I}(\bm{\theta}\,|\,\bm{x})$ (depending on both $\bm{\theta}$ and $\bm{x}$).
@@
In practice, the FIM is typically computationally expensive and needs to be estimated. Given $q(\bm{x})$ and a NN with weights and biases $\bm{\theta}$ parameterizing $p(\bm{y}\,|\,\bm{x};\bm{\theta})$, as per Eq. 1, we consider two commonly used estimators of the FIM Guo and Spall (2019); Soen and Sun (2021) given by
@@
^$\hat{\mathcal{I}}_{1}(\bm{\theta})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\frac{1}{N}\sum_{k=1}^{N}\left[\frac{\partial\log p(\bm{y}_{k}\,|\,\bm{x}_{k})}{\partial\bm{\theta}}\frac{\partial\log p(\bm{y}_{k}\,|\,\bm{x}_{k})}{\partial\bm{\theta}^{\top}}\right];\quad\textrm{and}\quad\hat{\mathcal{I}}_{2}(\bm{\theta})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\frac{1}{N}\sum_{k=1}^{N}\left[-\frac{\partial^{2}\log p(\bm{y}_{k}\,|\,\bm{x}_{k})}{\partial\bm{\theta}\partial\bm{\theta}^{\top}}\right],$ (3)
@@
where $p(\bm{y}_{k}\,|\,\bm{x}_{k})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}p(\bm{y}_{k}\,|\,\bm{x}_{k};\bm{\theta})$ and $(\bm{x}_{1},\bm{y}_{1}),\ldots,(\bm{x}_{N},\bm{y}_{N})$ are i.i.d. sampled from $p(\bm{x},\bm{y};\bm{\theta})$. A conditional variant of the estimators, denoted as ^$\hat{\mathcal{I}}_{1}(\bm{\theta}\,|\,\bm{x})$ and ^$\hat{\mathcal{I}}_{2}(\bm{\theta}\,|\,\bm{x})$, can be defined by fixing $\bm{x}=\bm{x}_{1}=\cdots=\bm{x}_{N}$ and sampling $\bm{y}_{1},\ldots,\bm{y}_{N}$ independently from $p(\bm{y}\,|\,\bm{x})$ in Eq. 3 — details omitted for brevity.
@@
Both estimators, ^$\hat{\mathcal{I}}_{1}(\bm{\theta})$ and ^$\hat{\mathcal{I}}_{2}(\bm{\theta})$, are random matrices with the same shape as $\mathcal{I}(\bm{\theta})$. By Eq. 2, they are unbiased — for ^$\hat{\mathcal{I}}_{2}(\bm{\theta})$, this only holds if activations functions are in $C^{2}(\Re)$. Following Eq. 1’s setting, the estimation variances of ^$\hat{\mathcal{I}}_{1}(\bm{\theta})$ and ^$\hat{\mathcal{I}}_{2}(\bm{\theta})$ can be expressed in closed form and upper bounded Soen and Sun (2021). This provides an important, yet not widely discussed, tool for quantifying the estimators’ accuracy Guo and Spall (2019) and hence insights for where / when different estimators should be used. Despite this, for deep NNs, neither these variances nor their bounds can be computed efficiently due to the huge dimensionality of $\bm{\theta}$.
This work focuses on estimating the diagonal entries of the FIM and their associated variances. Our results — including estimators of the FIM, their variances, and their variance bounds — can be implemented through automatic differentiation. These computational tools empower us to practically explore the trade-offs between the two estimators. For example, Fig. 1 shows natural gradient descent Amari (2016) for generalized linear models on a toy dataset, where ^$\hat{\mathcal{I}}_{2}(\bm{\theta})$ is preferable (especially for regression) and ^$\hat{\mathcal{I}}_{1}(\bm{\theta})$ suffers from high variance and an unstable learning curve. Our analytical results reveal how moments of the output exponential family and gradients of the NN in Eq. 1 affects the FIM estimators. We discover a general decomposition of the estimators’ variances corresponding to the samples of $\bm{x}$ and $\bm{y}$. We investigate different scenarios where each FIM estimator is the preferred one and then connect our analysis to the empirical FIM.
Figure 1: Natural gradient (NG) descent using ^$\hat{\mathcal{I}}_{1}(\bm{\theta})$ / ^$\hat{\mathcal{I}}_{2}(\bm{\theta})$ on a 2D toy dataset for regression (linear regression) and classification (logistic regression) (details in Appendix A). Inset plot shows the parameter updates throughout training. Here, the variance of ^$\hat{\mathcal{I}}_{2}(\bm{\theta})$ is generally lower than ^$\hat{\mathcal{I}}_{1}(\bm{\theta})$.
@@ 2 Related Work
Prior efforts aim to analyze the structure of the FIM of NNs with random weights Pennington and Worah (2018); Karakida et al. (2019; 2021); Amari et al. (2019); Papyan (2020). This body of work hinges on utilizing tools from random matrix theory and spectral analysis, characterizing the behavior and statistics of the FIM. One insight is that randomly weighted NNs have FIMs with a majority of eigenvalues close to zero; with the other eigenvalues taking large values Karakida et al. (2019; 2021). In our work, the randomness stems from sampling from data distributions $p(\bm{x},\bm{y})$ — which follows the principle of Monte Carlo (MC) information geometry Nielsen and Hadjeres (2019) that approximates information geometric quantities via MC estimation. We examine a different subject on how the distribution of the FIM on a matrix manifold is affected by finite sampling of the data distribution.
In the literature of NN optimization, a main focus is on deriving a computationally friendly proxy for the FIM. One can consider the unit-wise FIM Ollivier (2015); Le Roux et al. (2007); Sun and Nielsen (2017); Amari et al. (2019) (also known as quasi-diagonal FIM Ollivier (2015)), where a block-diagonal approximation of the FIM is taken to capture intra-neuron curvature information. Or one can consider the block-diagonal layer-wise FIM where each block corresponds to parameters within a layer  Kurita (1994); Martens et al. (2010); Pascanu and Bengio (2014); Martens and Grosse (2015); Heskes (2000); Ren and Goldfarb (2021); Karakida and Osawa (2020). NN optimizers can approximate the inverse FIM Singh and Alistarh (2020) or approximate the product of the inverse FIM and the gradient vector Ren and Goldfarb (2021). Much less attention is paid to how related approximations deviate from the true FIM Guo and Spall (2019); Soen and Sun (2021) or how optimization is affected by such deviation Sun and Spall (2021). For the univariate case, one can study the asymptotic variance of the Fisher information Guo and Spall (2019) with the central limit theorem. In deep NNs, the estimation variance of the FIM can be derived in closed form and bounded Soen and Sun (2021). However, our former analysis Soen and Sun (2021) has two limitations: (1) the variance tensors are 4D and can not be easily computed; (2) only the norm of these tensors are bounded, and it is not clear how the variance is distributed among individual parameters. The current work tackles these limitations by focusing on the diagonal elements of the FIM. Our results can be computed numerically at a reasonable cost in typical learning settings. We provide novel bounds so that one can quantify the accuracy of the FIM computation w.r.t. individual parameters or subgroup of parameters.
Issues of utilizing the empirical FIM to approximate the FIM have been highlighted Pascanu and Bengio (2014); Martens (2020). For example, estimators of the FIM do not in general capture any second-order information about the log-likelihood Kunstner et al. (2020). The empirical FIM is a biased estimator and can be connected with our unbiased estimators via a generalized definition of the Fisher matrix in Section 6.
Alternative to the FIM, the Generalized Gauss-Newton (GGN) matrix — a Hessian approximator — was originally motivated through the squared loss for non-linear models Martens (2020). The GGN is equivalent to the FIM when a loss function is taken to be the empirical expectation of the negative log-likelihood of Eq. 1 Heskes (2000); Pascanu and Bengio (2014); Martens (2020).
@@ 3 Variance of Diagonal FIM Estimators
In our notations, all vectors such as $\bm{x}$, $\bm{y}$, and $\bm{\theta}$ are column vectors. We use $k$ to index random samples $\bm{x}$ and $\bm{y}$ and use $i$ and $j$ to index the NN weights and biases $\bm{\theta}$. We shorthand $\bm{h}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\bm{h}_{\bm{\theta}}$, $p(\bm{y}\,|\,\bm{x})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}p(\bm{y}\,|\,\bm{x};\bm{\theta})$, and $p(\bm{y},\bm{x})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}p(\bm{y},\bm{x};\bm{\theta})$ whenever the parameters $\bm{\theta}$ is clear from context. To be consistent, we use ‘$\,|\,\bm{x}$ conditioning’ to distinguish between jointly calculated values versus conditioned values with fixed $\bm{x}$. By default, the derivatives are w.r.t. $\bm{\theta}$. For example, $\partial_{i}\bm{h}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\partial\bm{h}/\partial\theta_{i}$ and $\partial_{i}^{2}\bm{h}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\partial^{2}\bm{h}/\partial\theta_{i}^{2}$. We adopt Einstein notation to express tensor summations, so that an index appearing as both a subscript and a superscript in the same term indicates a summation. For example, $x^{a}y_{a}$ denotes $\sum_{a}x^{a}y_{a}$. For clarity, we mix standard $\Sigma$-sum and Einstein notation. We denote the variance and covariance of random variables by $\mathrm{Var}(\cdot)$ and $\mathrm{Cov}(\cdot)$, respectively.
@@
Based on the parametric form of the model in Eq. 1, the diagonal entries of the FIM estimators in Eq. 3 can be written as [^1] This and subsequent derivations can be found in the appendix.:
@@
^$\hat{\mathcal{I}}_{1}(\theta_{i})$ =.(^$\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\left(\hat{\mathcal{I}}_{1}(\bm{\theta})\right)_{ii}=\frac{1}{N}\sum_{k=1}^{N}\left(\frac{\partial F(\bm{h}(\bm{x}_{k}))}{\partial\theta_{i}}-\frac{\partial\bm{h}^{a}(\bm{x}_{k})}{\partial\theta_{i}}\cdot\bm{t}_{a}(\bm{y}_{k})\right)^{2};$